1) 99=3a
2) 6.5b=3
3) c/6=12
4) 4=d/12
5) c/0.3=150
6) 5=4f
7) g/12=16
8) 1.2h=3.6
9) 19=j/0.4
10) k/14=39

Answer by MArk_HeRras(6)

1) 99=3a
99=3a
(99)/(3)=(3a)/(3) First,Divide 3 to both sides
33=a
a=33 Thats The Answer
2) 6.5b=3
6.5b=3
(6.5b)/(6.5)=(3)/(6.5) First,Divide 6.5 to both sides
b=.46 The answer
3) c/6=12
c/6=12 First,Multiply 6 to both ides to get the answer
6*(c)/(6)=12*6
c=72 The answer
4) 4=d/12
4=d/12 The same as No.3.Multiply First, 12 to both sides
12*4=(d)/(12)* 12
48=d
d=48 The answer
5) c/0.3=150
c/0.3=150 The Same Also as No.3 and No.4.
.3*(c)/(0.3)=150*.3
c=5 The Answer
The Following 6-10 .I will answer it Nexttime.